OpenStudy (anonymous):

6(6X-1)-5(4X-18)=15X+15 x=? SIMPLIFY YOUR ANSWER

OpenStudy (lgbasallote):

distribute first then tell me what you get

OpenStudy (lgbasallote):

when i said tell me what you get i meant here not private messages hehe anyway i think you distributed wrong..let me try

OpenStudy (lgbasallote):

6(6x - 1) - 5(4x - 18) = 15x + 15 36x - 6 - 20x + 90 = 15x + 15 16x + 84 = 15x + 15 although i think you just made a typo

OpenStudy (lgbasallote):

anyway from here subtract 15x from both sides 16x - 15x + 84 = 15x - 15x + 15 x + 84 = 15 do you know what to do next?

OpenStudy (lgbasallote):

try subtracting 84 from both sides to make the 84 in the left side 0

OpenStudy (lgbasallote):

x + 84 - 84 = 15 - 84 x = 15 - 84

OpenStudy (lgbasallote):

what does this equal to